OpenStudy (anonymous):

Okay, I really need some help, can someone please check my answers!! I need it ASAP, so please help quick. I would really appreciate it!! :)

OpenStudy (anonymous):

1. What is the value of the y variable in the solution to the following system of equations? 3x + y = 11 –3x + 2y = 1 A. 2 B. 4 C. y can be any number as there are infinitely many solutions to the system. D. There is no y value as there is no solution to the system. 2. What is the value of the y variable in the solution to the following system of equations? 3x − 6y = 3 7x − 5y = −11 A. –2 B. –3 C. x can be any number as there are infinitely many solutions to the system. D. There is no x value as there is no solution to the system 3. What is the value of the x variable in the solution to the following system of equations? 4x + 2y = 12 x – y = 3 A. 0 B. –3 C. 3 D. 2 4. What is the value of the x variable in the solution to the following system of equations? 2x − y = 6 4x − 2y = 12 A. 0 B. 4 C. x can be any number as there are infinitely many solutions to the system. D. There is no x value as there is no solution to the system. 5. The three Math Idol judges have been eliminating contestants all day! The number of one-step equations and two-step equations who have been eliminated today is equal to 1120. If three times the number of one-step equations minus twice the number of two-step equations is equal to 1300, how many two-step equations auditioned today? A. 1300 B. 1120 C. 708 D. 412 My Answers: 1. B 2. A 3. C 4. D 5. D Am I right?

OpenStudy (anonymous):

I know it looks like a lot but its really only five questions and you would be just checking to see if im right... please someone help!!

OpenStudy (dumbcow):

looks good...the only one wrong is #4, it should be infinite solutions

OpenStudy (anonymous):

okay, thanks :)

OpenStudy (anonymous):

I really do appreciate it, so much :) :)

OpenStudy (dumbcow):

:)